OpenStudy (anonymous):

The parabola with vertex (2, 3), axis of symmetry parallel to the y-axis, and p = 4.

OpenStudy (anonymous):

\[ (y-y_0)= 4 p (x-x_0)^2 \]

OpenStudy (anonymous):

\[ y-3 = 16(x-2)^2\\ y = 3 + 16(x-2)^2\\ y=16 x^2-64 x+67 \]

OpenStudy (anonymous):

ok I have the formula but I am get confused on what is x and y and my answer supposed to end in a fraction equation

OpenStudy (anonymous):

Did you write the problem correctly?